OpenStudy (anonymous):

The value of the discriminant of the equation 4x^2+6x-3=0 is?

OpenStudy (anonymous):

The discriminant is the part of the quadratic formula under the square root sign, as that determines whether there is one solution (if it equals zero, as plus or minus zero is the same thing), two solutions (if it is positive), or two imaginary solutions (if it is negative, as the square root of a negative number is imaginary). It's the part that determines the nature of the solutions. Thus, we have:\[b^2-4ac=6^2-4(4)(-3)=36+48=84\]For more solutions to textbook problems, be sure to check out http://www.slader.com/s/eWFrZXlnbGVl. I submit a lot of content there as well.